# Improving Graph Neural Networks by Learning Continuous Edge Directions

### v1

@@ 5.1 Node classification
MLP 80.81$\pm$4.75 85.29$\pm$3.31 28.77$\pm$1.56 46.21$\pm$2.99 75.69$\pm$2.00
GCN 55.14$\pm$5.16 51.76$\pm$3.06 53.43$\pm$2.01 64.82$\pm$2.24 86.98$\pm$1.27
SAGE 82.43$\pm$6.14 81.18$\pm$5.56 41.61$\pm$0.74 58.73$\pm$1.68 86.90$\pm$1.04
GAT 52.16$\pm$6.63 49.41$\pm$4.09 40.72$\pm$1.55 60.26$\pm$2.50 86.33$\pm$0.48
GGCN 84.86$\pm$4.55 86.86$\pm$3.29 55.17$\pm$1.58 71.14$\pm$1.84 87.95$\pm$1.05
MagNet 83.3$\pm$6.1 85.7$\pm$3.2 39.01$\pm$1.93 58.22$\pm$2.87 82.63$\pm$1.80
FLODE 77.57$\pm$5.28 80.20$\pm$3.56 74.03$\pm$1.58 77.98$\pm$1.05 86.44$\pm$1.17
Sheaf 85.95$\pm$5.51 89.41$\pm$4.74 56.34$\pm$1.32 68.68$\pm$1.73 87.30$\pm$1.15
DirGNN 68.38$\pm$ 4.99 68.82$\pm$ 5.91 75.31$\pm$1.92 79.71$\pm$1.26 86.27$\pm$ 1.45
CoED 84.59$\pm$4.53 87.84$\pm$3.70 75.32$\pm$1.82 79.69$\pm$1.35 87.02$\pm$1.01
Table 1: Comparison of baseline models and CoED (without edge direction learning) for node classification task across different types of graphs. Top three models are colored by First, Second, Third. The reported numbers are the mean and standard deviation of test accuracies across different splits. The first two rows report the homophily ratios of the graphs and whether they are directed or undirected.
@@ 5.2 Node regression on graph ensemble dataset
Our key contribution is the joint learning of continuous edge directions alongside GNN parameters. This approach is most effective on graph ensemble data, where the graph structure remains fixed but multiple realizations of node features and targets exist. By learning edge directions for all edges without a need to mask parts of the graph, our method optimizes information flow across the entire graph. We empirically demonstrate the benefits of learning continuous edge directions using both synthetic and real-world graph ensemble datasets.
@@ 5.2.1 Synthetic datasets
Gene Regulatory Networks (GRNs) are directed graphs where nodes represent genes and edges represent interactions between pairs of genes (Figure 3(c)). In these networks, when two genes interact, one either activates or suppresses the other. We used Hill functions with randomly chosen parameters to define the dynamics of these gene-gene interactions. We constructed a directed GRN graph with 200 nodes and randomly assigned interactions between them. Starting from random initial expression levels, we solved the system of nonlinear ordinary differential equations representing the GRN dynamics to obtain the steady-state expression levels of all genes. Next, we modeled gene perturbations by setting the expression levels of either one or two genes (the perturbed set of genes) to zero and recomputing the steady-state expression levels for all genes using the same GRN dynamics (Figure 3(d)). We performed this procedure for all single-gene perturbations and a subset of double-gene perturbations, resulting in 1,200 different realizations. Our learning task is to predict the steady-state expression levels of all genes following perturbation (target node values) given the initial steady-state expression levels, with the perturbed genes set to zero (input node features). We provided baseline models with the original graph and CoED with the undirected version of the graph. Further details of the data generation are provided in Appendix A.2.2.
@@ Results.
As shown in Table 5, CoED achieves the best performance on both synthetic datasets by a considerable margin. For comparison, we include several baseline models: classical models, GCN and GAT; a transformer-based model with positional encoding, GraphGPS (Rampášek et al., 2022); a directionality-aware model, MagNet; a higher-order model, DRew (Gutteridge et al., 2023); and a combination of directionality-aware and higher-order model, FLODE. Details of the training setup, hyperparameter search procedure, and selected hyperparameters are provided in Appendix A.2.2.
To identify which aspects of GNNs are particularly effective for learning on graph ensemble datasets, we analyze the baseline models’ results in detail. On the undirected lattice graph, MagNet provides only a slight improvement over GCN, which is expected since MagNet reduces to ChebNet (Defferrard et al., 2016) on undirected graphs, and GCN is a first-order truncation of ChebNet. However, in the directed GRN experiments, MagNet shows substantial improvement over GCN. We also observe that higher-order GNNs like DRew and FLODE perform competitively on the undirected lattice graph. Notably, FLODE’s instantaneous enhancement of connectivity via fractional powers of the graph Laplacian outperforms DRew’s more gradual incorporation of higher-hop messages. However, both methods struggle on the directed GRN graph. On both synthetic graphs, attention-based models—GAT and GraphGPS—deliver strong performance, coming in just behind CoED. GraphGPS, in particular, seems to benefit from its final global attention step, similar to how FLODE benefits from densifying the graph. We also notice that increasing the dimension of Laplacian positional encoding does not further enhance GraphGPS’s performance. Interestingly, models that learn edge weights via attention mechanisms outperform MagNet on the directed GRN graph. This is likely because MagNet’s unitary evolution of complex-valued features does not resemble the actual feature propagation (i.e., the GRN dynamics), in addition to the shortcomings highlighted in the previous section.
We then investigated whether CoED can recover the true continuous edge directions of the triangular lattice graph, given that the feature propagation steps during data generation closely match the message-passing operation of CoED. As shown in Figure 5, CoED correctly learns the true directions. Lastly, since both synthetic datasets are generated by propagating input features over multiple hops, we investigated how performance scales with model depth by training CoED and the second-best model with up to 10 layers. Figure 6 demonstrates that CoED continues to improve as depth increases, while the performance of the other models plateau at a shallower depth.
@@ 5.2.2 Real Datasets
Perturb-seq (Dixit et al., 2016) is a well-established experimental technique in single-cell biology that inspired the synthetic GRN experiment described earlier. In Perturb-seq experiments, one or more genes in a cell are knocked out resulting in zero expresion—as in our synthetic GRN dataset. The resulting changes in the expression levels of all other genes are then measured to elucidate gene-gene interactions. For our study, we used the Replogle-gwps dataset (Replogle et al., 2022; Peidli et al., 2024), which includes 9,867 distinct single-gene perturbations, along with control measurements from cells without any perturbation to establish baseline gene expression levels. The learning task is again predicting the expression levels of all genes following perturbation given the initial steady-state with the expression levels of the perturbed genes set to zero. Since there is no ground truth gene regulatory network (GRN) available for this dataset, we constructed an undirected $k$-nearest neighbors graph to connect genes with highly correlated expression patterns. All models are trained using this heuristic graph. Details of the data processing procedure are provided in Appendix A.3.1.
@@
We also modeled the traffic flow between Wikipedia pages using the WikiMath dataset, which is classified as a “static graph with temporal signals” in the PyTorch Geometric Temporal library (Rozemberczki et al., 2021). In this dataset, each node corresponds to an article page on a popular mathematics topic, and each directed edge represents a link from one page to another. The node features are the daily visit counts of all pages over a period of 731 consecutive days. The learning task is node regression: predict the next day’s visit counts across all pages given today’s visit counts. We trained the baseline models using the true, directed graph, while CoED was trained starting from the undirected version of the graph. Additional details are provided in Appendix A.3.2.
@@ Results.
Table 2 reports the test performances of all baseline models and CoED on the three datasets. The baselines include GAT, MagNet, DirGCN, and DirGAT. We focus on these models because attention-based approaches showed competitive performance on the synthetic datasets, and MagNet, which accounts for edge directions, performed well on the directed GRN dataset. Details of the training setup, hyperparameter search procedure, and selected hyperparameters are provided in the Appendix.
We observe that CoED achieves the best performance across all three datasets. On the Perturb-seq dataset with an undirected graph, MagNet performs similarly to GCN, while DirGCN struggles. We attribute DirGCN’s poor performance to clashing learnable parameters: it uses two distinct weight matrices, $\mathbf{W}_{\leftarrow}$ and $\mathbf{W}_{\rightarrow}$, applied to identical in- and out-neighbor aggregated messages in the case of undirected graphs. For a propagation path of $L$ hops, this results in $2^{L}$ feature transformations, comprised of different combinations of the two weight matrices, which together reduce the model’s ability to efficiently learn the optimal weight matrices. In contrast, DirGAT’s attention mechanisms break the symmetry of the undirected edges, leading to improved performance. CoED naturally addresses this issue in undirected graphs by learning the edge directions. On the web traffic and power grid datasets, which have directed graphs, we observe a similar trend. MagNet outperforms GCN due to its ability to process directed messages. However, GAT delivers better performance than MagNet, likely because its attention mechanism effectively captures important features. Since directed graphs create distinct feature propagation paths, DirGCN achieves substantial performance gains. DirGAT further improves upon DirGCN by leveraging additional edge weight learning through an attention mechanism. CoED surpasses all these models, demonstrating the effectiveness of learning continuous edge directions.
@@ 6 Conclusion
We have introduced the Continuous Edge Direction (CoED) GNN, which assigns fuzzy, continuous directions to the edges of a graph and employs a novel complex-valued Laplacian to transform information propagation on graphs from diffusion to directional flow. Our theoretical analysis shows that CoED GNN is more expressive than existing Laplacian-based methods and matches the expressiveness of an extended Weisfeiler-Leman test for directed graphs with fuzzy edges. Through extensive experiments on both synthetic and real-world graph ensemble datasets—including gene regulatory networks, web traffic, and power grids—we demonstrated that learning continuous edge directions significantly improves performance over existing GNN models. These results highlight CoED GNN’s effectiveness in enabling long-range information flow, offering a powerful framework for learning on graphs.
@@ A.1 Node classification
All the results reported in Table 1 are taken from (Bodnar et al., 2022) except for MagNet, FLODE, and DirGNN. Magnet’s results on Texas and Wisconsin are taken from the original paper, while those on Squirrel and Chameleon are taken from (Rossi et al., 2024). FLODE’s results on Squirrel, Chameleon, and Cora, and DirGNN’s results on Squirrel and Chameleon are taken from their respective papers. We trained these models and CoED for the remaining results and describe the training procedures below. All five datasets were downloaded using PyTorch Geometric library (Fey & Lenssen, 2019) with split=geom-gcn’ argument to use the 10 fixed 48%/32%/20% training/validation/test splits provided by (Pei et al., 2020), which are the splits used in all the referenced results.
@@
# layers # hidden lr wd dropout self-loop K q $\alpha$ norm jk self-feature
MagNet (Cora) 1 32 1e-2 0 0.5 1 2 0 - - - -
FLODE (Texas) 1 128 1e-2 1e-3 0 0 - - - - - -
FLODE (Wisconsin) 1 128 5e-3 0 0.5 0 - - - - - -
DirGNN (Cora) 2 64 5e-3 5e-4 0.5 - - - 0 False None -
DirGNN (Texas) 4 64 5e-3 1e-3 0.5 - - - 0.5 False None -
DirGNN (Wisconsin) 2 128 5e-3 1e-4 0.5 - - - 0.5 False None -
CoED (Texas) 2 64 2e-2 5e-4 0.5 0 - - 0.5 False None True
CoED (Wisconsin) 2 128 2e-2 1e-3 0.5 0 - - 0.5 False None True
CoED (Squirrel) 3 128 1e-2 0 0 0 - - 0 True max False
CoED (Chameleon) 4 128 5e-3 0 0 0 - - 0 True cat False
CoED (Cora) 2 128 5e-4 1e-4 0.5 1 - - 0 False None False
Table A.1: Hyperparameters selected for node classification experiments

### v2

@@ 5.1 Node classification without edge direction learning
MLP 64.94$\pm$0.62 31.34$\pm$0.05 80.81$\pm$4.75 85.29$\pm$3.31 36.70$\pm$0.21 37.53$\pm$1.74 39.05$\pm$3.74 74.02$\pm$1.90 83.56$\pm$0.26 90.75$\pm$0.31 75.69$\pm$2.00
GCN 73.69$\pm$0.74 51.02$\pm$0.06 55.14$\pm$5.16 51.76$\pm$3.06 46.02$\pm$0.26 39.47$\pm$1.47 40.89$\pm$4.12 76.50$\pm$1.36 89.65$\pm$0.52 92.70$\pm$0.20 86.98$\pm$1.27
SAGE 85.74$\pm$0.67 48.43$\pm$0.21 82.43$\pm$6.14 81.18$\pm$5.56 52.94$\pm$0.14 36.09$\pm$1.99 37.77$\pm$4.14 76.04$\pm$1.30 91.20$\pm$0.29 94.59$\pm$0.14 86.90$\pm$1.04
GAT 80.87$\pm$0.30 45.92$\pm$0.22 52.16$\pm$6.63 49.41$\pm$4.09 46.05$\pm$0.51 35.62$\pm$2.06 39.21$\pm$3.08 76.55$\pm$1.23 90.78$\pm$0.13 93.87$\pm$0.11 86.33$\pm$0.48
GGCN 74.46$\pm$0.54 OOM 84.86$\pm$4.55 86.86$\pm$3.29 OOM 37.46$\pm$1.57 38.71$\pm$3.04 77.14$\pm$1.45 91.81$\pm$0.20 94.50$\pm$0.11 87.95$\pm$1.05
FLODE 74.97$\pm$0.53 OOM 77.57$\pm$5.28 80.20$\pm$3.56 OOM 38.63$\pm$1.68 42.85$\pm$3.89 78.07 $\pm$1.62 90.88$\pm$0.23 95.93$\pm$0.20 86.44$\pm$1.17
Sheaf 77.94$\pm$0.53 OOM 85.95$\pm$5.51 89.41$\pm$4.74 48.77$\pm$0.20 39.03$\pm$1.73 41.98$\pm$3.42 77.14$\pm$1.85 90.56$\pm$0.13 95.01$\pm$ 0.17 87.30$\pm$1.15
MagNet 88.07$\pm$0.27 OOM 83.3$\pm$6.1 85.7$\pm$3.2 60.29$\pm$0.27 42.7$\pm$1.5 44.5$\pm$1.1 75.26$\pm$1.63 90.30$\pm$0.27 94.54$\pm$0.19 82.63$\pm$1.80
Chung 87.35$\pm$0.53 64.77$\pm$0.23 80.54$\pm$4.65 81.79$\pm$5.42 53.01$\pm$0.45 42.46$\pm$1.77 43.47$\pm$3.64 76.08$\pm$1.11 92.57$\pm$0.16 95.47$\pm$0.14 86.03$\pm$1.63
DirGNN 91.23$\pm$ 0.32 73.95$\pm$0.05 83.78 $\pm$ 2.70 85.88$\pm$2.11 64.08$\pm$0.26 44.19$\pm$2.42 46.08$\pm$ 2.67 76.63$\pm$1.51 92.97$\pm$0.26 96.13$\pm$0.12 86.27$\pm$ 1.45
Co-GNN 91.57$\pm$0.32 48.31$\pm$0.15 83.51$\pm$5.19 86.47$\pm$3.77 49.82$\pm$0.24 39.85$\pm$1.15 41.92$\pm$4.03 76.49$\pm$1.40 92.76$\pm$0.22 95.95$\pm$0.14 87.44$\pm$ 0.85
CoED 92.17$\pm$0.29 74.67$\pm$0.02 84.59$\pm$4.53 87.84$\pm$3.70 64.59$\pm$0.20 45.50$\pm$1.62 47.27$\pm$3.62 77.14$\pm$1.57 92.88$\pm$0.15 95.83$\pm$0.12 87.02$\pm$1.01
Table 1: Comparison of baseline models and CoED (without edge direction learning) for node classification task across different types of graphs. Top three models are colored by First, Second, Third. The reported numbers are the mean and standard deviation of test accuracies across different splits. The first two rows report the homophily ratios of the graphs and whether they are directed or undirected. OOM indicates out-of-memory error.
@@ 5.2 Node regression on graph ensemble dataset
Our key contribution is the joint learning of continuous edge directions alongside GNN parameters. This approach is particularly effective on graph ensemble data, where the graph structure remains fixed but multiple realizations of node features and targets exist. By learning edge directions for all edges without a need to mask parts of the graph, our method optimizes information flow across the entire graph. Learning edge directions for the node classification task above, where a subset of nodes are masked for testing, would optimize the edges connected to the training node at the expense of those connected to the test nodes, diminishing overall performance. However, in the graph ensemble setting, learning continuous edge directions substantially improves performance, as we empirically demonstrate below on both synthetic and real-world datasets.
@@ 5.2.1 Synthetic datasets
Gene Regulatory Networks (GRNs) are directed graphs where nodes represent genes and edges represent interactions between pairs of genes (Figure 3(c)). In these networks, when two genes interact, one either activates or suppresses the other. We used Hill functions with randomly chosen parameters to define the dynamics of these gene-gene interactions. We constructed a directed GRN graph with 200 nodes and randomly assigned interactions between them. Starting from random initial expression levels, we solved the system of nonlinear ordinary differential equations representing the GRN dynamics to obtain the steady-state expression levels of all genes. Next, we modeled gene perturbations by setting the expression levels of either one or two genes (the perturbed set of genes) to zero and recomputing the steady-state expression levels for all genes using the same GRN dynamics (Figure 3(d)). We performed this procedure for all single-gene perturbations and a subset of double-gene perturbations, resulting in 1,200 different realizations. Our learning task is to predict the steady-state expression levels of all genes following perturbation (target node values) given the initial steady-state expression levels with the perturbed genes set to zero (input node features). We provided baseline models with the original graph and CoED with the undirected version of the graph. Further details of the data generation are provided in Appendix A.2.2.
@@ Results.
Table 5 shows the test performances of CoED alongside several baseline models: classical models, GCN and GAT; a transformer-based model with positional encoding, GraphGPS (Rampášek et al., 2022); a directionality-aware model, MagNet and the model based on the Chung Laplacian; a higher-order model, DRew (Gutteridge et al., 2023); and a combination of directionality-aware and higher-order model, FLODE. Details of the training setup, hyperparameter search procedure, and selected hyperparameters are provided in Appendix A.2.2.
To identify which aspects of GNNs are particularly effective for learning on graph ensemble datasets, we analyze the baseline models’ results in detail. On the undirected lattice graph, MagNet provides only a slight improvement over GCN, which is expected since MagNet reduces to ChebNet (Defferrard et al., 2016) on undirected graphs, and GCN is a first-order truncation of ChebNet. However, in the directed GRN experiments, MagNet shows substantial improvement over GCN. We also observe that higher-order GNNs like DRew and FLODE perform competitively on the undirected lattice graph. Notably, FLODE’s instantaneous enhancement of connectivity via fractional powers of the graph Laplacian outperforms DRew’s more gradual incorporation of higher-hop messages. However, both methods struggle on the directed GRN graph. The model based on the Chung Laplacian demonstrates strong performance on the lattice graph but offers only a modest improvement over GCN on the directed GRN. On both synthetic graphs, attention-based models—GAT and GraphGPS—deliver strong performance, coming in just behind CoED. GraphGPS, in particular, seems to benefit from its final global attention step, similar to how FLODE benefits from densifying the graph. We also notice that increasing the dimension of Laplacian positional encoding does not further enhance GraphGPS’s performance. Interestingly, models that learn edge weights via attention mechanisms outperform MagNet on the directed GRN graph. This is likely because MagNet’s unitary evolution of complex-valued features does not resemble the actual feature propagation (i.e., the GRN dynamics), in addition to the shortcomings highlighted in the previous section. CoED outperforms the attention-based models in both datasets as it optimizes the edge directions directly as a part of the learning objective independent of the node features. The shortcomings of GAT compared to CoED on graph ensemble data are discussed further in Appendix H.
We then investigated whether CoED can recover the ground truth continuous edge directions of the triangular lattice graph, given that the feature propagation steps during data generation closely resemble the message-passing operation of CoED. As shown in Figure 5, CoED correctly learns the true directions. Lastly, since both synthetic datasets are generated by propagating input features over multiple hops, we investigated how performance scales with model depth by training CoED and the second-best model with up to 10 layers. Figure 6 demonstrates that CoED continues to improve as depth increases, while the performance of the other models plateau at a shallower depth.
@@ 5.2.2 Real Datasets
Perturb-seq (Dixit et al., 2016) is a well-established experimental technique in single-cell biology that inspired the synthetic GRN experiment described earlier. In Perturb-seq experiments, one or more genes in a cell are knocked out resulting in zero expression—as in our synthetic GRN dataset. The resulting changes in the expression levels of all other genes are then measured to elucidate gene-gene interactions. For our study, we used the Replogle-gwps dataset (Replogle et al., 2022; Peidli et al., 2024), which includes 9,867 distinct single-gene perturbations, along with control measurements from cells without any perturbation to establish baseline gene expression levels. The learning task is again predicting the expression levels of all genes following perturbation given the initial steady-state with the expression levels of the perturbed genes set to zero. Since there is no ground truth gene regulatory network (GRN) available for this dataset, we constructed an undirected $k$-nearest neighbors graph to connect genes with highly correlated expression levels. All models are trained using this heuristic graph. Details of the data processing procedure are provided in Appendix A.3.1.
@@
We also modeled the traffic flow between Wikipedia articles using the WikiMath dataset, which is classified as a “static graph with temporal signals” in the PyTorch Geometric Temporal library (Rozemberczki et al., 2021). In this dataset, each node corresponds to a Wikipedia article on a popular mathematics topic, and each directed edge represents a link from one article to another. The node features are the daily visit counts of all articles over a period of 731 consecutive days. The learning task is node regression: predict the next day’s visit counts across all articles given today’s visit counts. We trained the baseline models using the ground truth directed graph, while CoED was trained starting from the undirected version of the graph. Additional details are provided in Appendix A.3.2.
@@ Results.
Table 2 reports the test performances of all baseline models and CoED on the three datasets. The baselines include GAT, MagNet, DirGCN, and DirGAT. We focus on these models because attention-based approaches showed competitive performance on the synthetic datasets, and MagNet, which accounts for edge directions, performed well on the directed GRN dataset. Details of the training setup, hyperparameter search procedure, and selected hyperparameters are provided in Appendix A.3.
We observe that CoED achieves the best performance across all three datasets. On the Perturb-seq dataset with an undirected graph, MagNet performs similarly to GCN while DirGCN struggles. We attribute DirGCN’s poor performance to clashing learnable parameters: it uses two distinct weight matrices, $\mathbf{W}_{\leftarrow}$ and $\mathbf{W}_{\rightarrow}$, applied to identical in- and out-neighbor aggregated messages in the case of undirected graphs. For a propagation path of $L$ hops, this results in $2^{L}$ feature transformations, comprised of different combinations of the two weight matrices, which together reduce the model’s ability to efficiently learn the optimal weight matrices. In contrast, DirGAT’s attention mechanisms break the symmetry of the undirected edges, leading to improved performance. CoED naturally addresses this issue by learning the edge directions, which are visualized in Appendix I.1. On the web traffic and power grid datasets, which have directed graphs, we observe a similar trend. MagNet outperforms GCN due to its ability to process directed messages. However, GAT delivers better performance than MagNet, likely because its attention mechanism effectively captures important features. Since directed graphs create distinct feature propagation paths, DirGCN achieves substantial performance gains. DirGAT further improves upon DirGCN by leveraging additional edge weight learning through an attention mechanism. CoED surpasses all these models, demonstrating the effectiveness of learning continuous edge directions.
@@ 6 Conclusion
We have introduced the Continuous Edge Direction (CoED) GNN, which assigns fuzzy, continuous directions to the edges of a graph and employs a novel complex-valued Laplacian to transform information propagation on graphs from diffusion to directional flow. Our theoretical analysis shows that CoED GNN is more expressive than existing Laplacian-based methods and matches the expressiveness of an extended Weisfeiler-Leman (WL) test for directed graphs with fuzzy edges. Through extensive experiments on both synthetic and real-world graph ensemble datasets—including gene regulatory networks, web traffic, and power grids—we demonstrated that learning continuous edge directions significantly improves performance over existing GNN models.
Acknowledgement
The authors thank Shishir Adhikari for his contributions to the conceptual formulation of this work and for numerous insightful discussions on graph Laplacians on directed graphs. We acknowledge funding from the National Institutes of Health (NIH)–National Heart, Lung, and Blood Institute (NHLBI)–under grant R01HL158269, and Advanced Research Projects Agency for Health (ARPA-H) award 1AYSAX000005.
@@ A.1 Node classification
In Table 1, MLP, GCN, SAGE, GAT, GGCN, and Sheaf’s results on Texas, Wisconsin, Citeseer, and Cora are taken from Bodnar et al. (2022); MLP, GCN, MagNet, and DirGNN’s results on Roman-Empire, SNAP-Patents, and Arxiv-Year from Rossi et al. (2024); SAGE and GAT’s results on Roman-Empire, and GCN, SAGE, and GAT’s results on the filtered versions of Squirrel and Chameleon from Platonov et al. (2023); GCN, SAGE, GAT, and GGCN’s results on AM-Computers and AM-Photo, and GGCN’s result on Roman-Empire from Deng et al. (2024), SAGE and GAT’s results on SNAP-Patents from Dwivedi et al. (2023); MagNet’s result on Texas and Wisconsin from the original paper (Zhang et al., 2021) and the filtered versions of Squirrel and Chameleon from (Sun et al., 2024); Flode’s results on Roman-Empire, Citeseer, and Cora from the original paper (Maskey et al., 2023); Co-GNN’s results on Roman-Empire and Cora from the original paper (Finkelshtein et al., 2023); GAT’s result on Arxiv-Year from Lim et al. (2021). We trained CoED and baseline models to fill the remaining entries in the table. We describe the training procedures below. Texas, Wisconsin, Citeseer, and Cora datasets were downloaded using PyTorch Geometric library (Fey & Lenssen, 2019) with split=geom-gcn’ argument to use the 10 fixed 48%/32%/20% training/validation/test splits provided by (Pei et al., 2020). We downloaded AM-Computers and AM-Photo datasets from the same library but used the 60%/20%/20%-split file provided in the repository of Deng et al. (2024). Since the original Squirrel and Chameleon datasets (Pei et al., 2020) have redundant nodes, we used the filtered versions with directed graphs provided in the repository of Platonov et al. (2023). For Roman-Empire, SNAP-Patents, and Arxiv-Year datasets, we used the dataloading pipeline provided in the repository of Rossi et al. (2024). All referenced results used the same splits as in our experiments.
@@
# layers # hidden lr wd dropout self-loop $\alpha$ norm jk self-feature
Roman-Empire 5 256 1e-2 0 0.2 0 0.5 False max True
SNAP-Patents 5 32 1e-2 0 0 0 0.5 True cat True
Texas 2 64 2e-2 5e-4 0.5 0 0.5 False None True
Wisconsin 2 128 2e-2 1e-3 0.5 0 0.5 False None True
Arxiv-Year 6 256 5e-3 0 0 0 0.5 False max False
Squirrel 3 64 1e-2 5e-3 0 0 0.5 False cat False
Chameleon 2 64 1e-2 2e-3 0 0 0.5 False cat False
Citeseer 2 256 2e-3 0 0.7 0 0.5 False None True
AM-Computers 2 512 5e-3 0 0.7 1 0 False None False
AM-Photo 3 128 1e-3 0 0.7 1 0.5 False None True
Cora 2 128 5e-4 1e-4 0.5 1 0 False None False
Table A.1: Selected hyperparameters CoED.
# env/act layers # env/act hidden $\tau_{0}$ conv type lr wd dropout layer norm skip
SNAP-Patents 2 / 1 32 / 4 0 GCN 1e-2 0 0 True False
Texas 3 / 1 64 / 4 0.1 GCN 2e-2 1e-3 0.5 False True
Wisconsin 4 / 2 64 / 4 0.1 GCN 2e-2 5e-4 0.5 False True
Arxiv-Year 3 / 1 128 / 8 0.1 GCN 1e-3 0 0 True True
Squirrel 2 / 1 64 / 4 0.1 Mean GCN 5e-3 0 0 False False
Chameleon 2 / 1 64 / 4 0.1 Mean GCN 5e-3 0 0 False False
Citeseer 4 / 1 64 / 4 0 GCN 1e-2 0 0.7 True True
AM-Computers 4 / 2 64 / 4 0 GCN 5e-3 0 0 True False
AM-Photo 3 / 2 64 / 8 0 GCN 5e-3 0 0.5 True True
Table A.2: Selected Hyperparameters for Co-GNN.
# layers # hidden conv type lr wd dropout $\alpha$ norm jk
Texas 2 256 DirSAGE 2e-2 5e-4 0.5 0.5 False None
Wisconsin 3 256 DirSAGE 1e-2 1e-4 0.5 0.5 False None
Squirrel 4 64 DirGCN 1e-2 5e-3 0 0.5 False cat
Chameleon 2 64 DirGCN 1e-2 2e-3 0 0.5 False cat
Citeseer 2 128 DirSAGE 5e-3 1e-3 0.5 0.5 False None
AM-Computers 3 256 DirSAGE 1e-3 0 0.5 0.5 True None
AM-Photo 4 128 DirSAGE 1e-3 0 0.5 0.5 True None
Cora 2 64 DirGCN 5e-3 5e-4 0.5 0 False None
Table A.3: Selected hyperparameters for DirGNN.
# layers # hidden lr wd dropout $\alpha$ norm jk self-feature
Roman-Empire 5 256 1e-2 0 0.2 0.5 False cat True
SNAP-Patents 5 32 1e-2 0 0 0.5 True cat True
Texas 3 64 1e-2 1e-3 0 0.5 True None True
Wisconsin 2 32 1e-2 1e-4 0.5 0.5 False None True
Arxiv-Year 6 256 5e-3 0 0 0.5 False cat False
Squirrel 2 32 5e-3 5e-3 0 0.5 False cat False
Chameleon 2 128 2e-2 1e-3 0 0.5 False cat False
Citeseer 2 64 1e-3 1e-4 0.5 0 False None True
AM-Computers 3 256 1e-3 0 0 0 True None True
AM-Photo 4 256 5e-3 1e-4 0 0 True None True
Cora 2 32 5e-4 0 0.5 0 False None True
Table A.4: Selected hyperparameters for Chung.
# layers # hidden lr wd dropout K q
Citeseer 3 256 1e-3 0 0.5 2 0
AM-Computers 5 256 1e-3 1e-4 0.5 2 0
AM-Photo 5 256 2e-3 0 0.7 3 0
Cora 1 32 1e-2 0 0.5 2 0
Table A.5: Selected hyperparameters for MagNet.
# layers # hidden type # stalk lr wd sheaf wd dropout high/low-pass filter # encoder layers
Roman-Empire 6 16 Diagonal 3 2e-3 1e-2 0 0.7 False/True 2
Arxiv-Year 8 16 Diagonal 4 1e-2 0 0 0 True/False 1
Squirrel 2 8 Orthogonal 6 1e-3 0 0 0 Ture/False 1
Chameleon 2 16 Diagonal 6 1e-2 0 0 0 True/False 1
AM-Computers 4 32 General 6 1e-2 0 0 0 True/False 1
AM-Photo 4 16 Diagonal 8 1e-2 0 0 0 True/False 1
Table A.6: Selected hyperparameters for Sheaf.
# layers # hidden # encoder/decoder layers lr wd dropout self-loop
Texas 1 128 1 / 1 1e-2 0 0 0
Wisconsin 1 128 1 / 1 5e-3 0 0.5 0
Squirrel 1 16 1 / 1 2e-2 1e-2 0.5 1
Chameleon 1 256 1 / 1 1e-2 5e-3 0.5 1
AM-Computers 1 32 1 / 2 5e-3 0 0 1
AM-Photo 1 64 2 / 1 5e-3 5e-3 0.5 1
Table A.7: Selected hyperparameters for FLODE.
# layers # hidden lr wd dropout
MLP (Squirrel) 3 64 1e-3 1e-4 0.5
MLP (Chameleon) 4 128 1e-2 0 0.7
MLP (AM-Computers) 2 256 5e-3 1e-4 0
MLP (AM-Photo) 2 256 5e-3 1e-4 0
SAGE (Arxiv-Year) 4 128 1e-3 1e-4 0
GGCN (Squirrel) 2 32 5e-3 1e-4 0.7
GGCN (Chameleon) 4 64 1e-2 5e-3 0.7
Table A.8: Selected hyperparameters for MLP, SAGE, and GGCN.